OpenStudy (anonymous):

hey :) Identify the vertex, focus, and directrix of the parabola with the equation x^2 – 6x – 8y + 49 = 0

OpenStudy (anonymous):

@experimentX hey :) i'm new to this...but do you know how to do this??

OpenStudy (experimentx):

do you know the standard equation of Parabola?

OpenStudy (anonymous):

ughh, is it ax^2+bx+c

OpenStudy (experimentx):

no google it

OpenStudy (anonymous):

y - k = a(x - h)^2?

OpenStudy (experimentx):

yeah ... can you find directrix, focus and vertex from it?

OpenStudy (anonymous):

hm, i don't think so..i get so confused on this stuff :(

OpenStudy (experimentx):

do you know what these terms mean?

OpenStudy (anonymous):

yeah, sort of

OpenStudy (anonymous):

directrix is a line that the parabola that curves away from

OpenStudy (experimentx):

not just that ...

OpenStudy (anonymous):

i forgot :/ but i know first with have to put the equation into standard form (y - k)² = 4p(x - h) ....sooo

OpenStudy (experimentx):

looks like we have excellent video on khan academy http://www.youtube.com/watch?v=ZJf9shWlMz0

OpenStudy (anonymous):

it's 14 minutes longg, but okay :)

OpenStudy (experimentx):

also here http://www.youtube.com/watch?v=ZJf9shWlMz0

OpenStudy (anonymous):

ooh, i remember this stuff now! kindaa :)

OpenStudy (anonymous):

but how do i put 6x – 8y + 49 = 0 into like standard form equation

OpenStudy (anonymous):

x2 – 6x – 8y + 49 = 0 **

OpenStudy (anonymous):

i don't have that much time to do this :(

OpenStudy (experimentx):

no ...